OpenStudy (mariebrittany07):

how would you figure out 2/3x+4=3/5x-2

OpenStudy (anonymous):

2/3x+4=3/5x-2 \[\frac{ 2 }{ 3x }+4 = \frac{ 3 }{ 5x }-2\]is it